OpenStudy (anonymous):

(x+3)/(2x) * (4x^2)/(2x-1)

OpenStudy (agent0smith):

\[\Large \frac {(x+3)} {2x} * \frac {4x^2} {(2x-1)}= \] change the order to make it easier\[\Large \frac {(x+3)} {2x-1} * \frac {4x^2} {2x}\]

OpenStudy (agent0smith):

@Katelynmorris ...?

OpenStudy (anonymous):

Not quite sure how to work out the problem from there?

OpenStudy (anonymous):

Does the (4x^2)/(2x) equal 2x^2? @agent0smith

OpenStudy (agent0smith):

what is x^2/x? (it's not x^2) \[\Large \frac{ x^a }{ x^b}= x^{a-b}\]

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Yes. \[\Large \frac {(x+3)} {2x-1} * 2x\]

OpenStudy (anonymous):

(2x^2+6)/(2x-1)?

OpenStudy (anonymous):

6x*

OpenStudy (agent0smith):

yep. \[\Large \frac {2x(x+3)} {2x-1} =\frac {2x^2+6x} {2x-1}\]i think either answer is fine though

OpenStudy (anonymous):

Thank you!